# Language Lives in Sparse Dimensions: Toward Interpretable and Efficient Multilingual Control for Large Language Models

### v1

@@ Abstract
Large language models exhibit strong multilingual capabilities despite limited exposure to non-English data. Prior studies show that English-centric large language models map multilingual content into English-aligned representations at intermediate layers and then project them back into target-language token spaces in the final layer. From this observation, we hypothesize that this cross-lingual transition is governed by a small and sparse set of dimensions, which occur at consistent indices across the intermediate to final layers. Building on this insight, we introduce a simple, training-free method to identify and manipulate these dimensions, requiring only as few as 50 sentences of either parallel or monolingual data. Experiments on a multilingual generation control task reveal the interpretability of these dimensions, demonstrating that the interventions in these dimensions can switch the output language while preserving semantic content, and that it surpasses the performance of prior neuron-based approaches at a substantially lower cost. We will release the code soon. [^1] https://github.com/ku-nlp/language-specific-dimensions
@@
Chengzhi Zhong1,   Fei Cheng1,    Qianying Liu2, Yugo Murawaki1,Chenhui Chu1,Sadao Kurohashi1,2 1 Kyoto University, Japan 2 National Institute of Informatics, Japan {zhong, feicheng, murawaki, chu, kuro}@nlp.ist.i.kyoto-u.ac.jp ying@nii.ac.jp
@@ 4.2 Data for Identification
4.3 Multilingual Generation Control
We evaluate with a pipeline of three metrics:
  • BLEU: evaluates translation quality for successful samples.
@@ 4.3 Multilingual Generation Control
Llama2-7b Neuron Kojima 42.0 15.3 55.7 15.1 81.3 10.2 57.5 7.5 77.0 16.6 60.8 12.8 7.76
Neuron Tang 83.7 28.2 72.7 21.1 86.7 17.9 20.5 14.4 98.0 20.8 72.3 20.9 15.14
@@
Llama2-13b Neuron Kojima 15.7 7.3 30.3 9.5 96.0 15.9 64.5 8.6 14.0 4.3 47.4 9.9 4.71
Neuron Tang 83.7 32.5 14.7 23.5 99.3 19.1 64.5 8.6 42.0 25.8 53.9 21.4 11.53
@@
Llama3.1-8b Neuron Kojima - -
Neuron Tang - -
Ours (Mono) 97.2 28.4 97.6 19.0 93.9 21.0 80.8 11.3 99.0 18.4 93.9 20.7 19.47
@@
Aya23-8b Neuron Kojima - -
Neuron Tang - -
Ours (Monolingual) 68.0 21.7 69.4 14.7 22.8 20.3 69.8 10.8 58.7 14.0 55.6 16.5 9.34
@@
5 Results and Discussion
We first search for an appropriate top-$K$ threshold for selecting language-specific dimensions. Therefore, we run the multilingual generation control experiment on Llama2-7B, applying an intervention at layer 20 with $\alpha\!=\!0.4$ in the parallel setting while varying $K$. As shown in Fig. 3, as $K$ increases, ACC improves but BLEU drops. The composite ACC*BLEU improves from 13.62 at K=50K\!=\!50, to 14.91 at K=100K\!=\!100 (+1.29), and to 15.80 at K=400K\!=\!400 (+2.18). Considering this trade-off, the most essential language-specific dimensions appear to be captured at around K=100K\!=\!100, which yields a smaller loss in semantic fidelity while still improving accuracy. However, to maximize performance, we set K=400K\!=\!400 for the subsequent experiments, which is about 7.8% of the hidden size for Llama2-13B and 9.8% for the other models.
In the monolingual setting, we need to choose an intermediate reference layer to compare with the final layer. We therefore set K=400K\!=\!400, apply an intervention at layer 20 with $\alpha\!=\!0.4$, and evaluate Llama2-7B on the multilingual generation control task while varying the layer $l$ from which dimensions are identified. As shown in Fig. 4, ACC and BLEU remain stable over a broad range of $l$, and the composite ACC*BLEU attains its best value near l=20l\!=\!20 but degrades at very deep layers. This indicates that the results are not highly sensitive to the choice of intermediate layer, and we use l=20l\!=\!20 in subsequent experiments.
@@ 5.2 Main Results of Controlling Multilingual Generation
To assess generality, we evaluate on two latest models—Llama3.1–8B and Aya23–8B. On Llama3.1–8B in the monolingual setting, the overall ACC*BLEU reaches 18.70, the best among the settings we evaluated. On Aya23–8B in the parallel setting, it reaches 10.51. These results support our hypothesis about language-specific dimensions and indicate broad applicability across model families. We omit neuron-based baselines, as they are not compatible with these newer models.
5.3 Are 50 Sentences Sufficient for Identification
@@ 6 Conclusion
In this work, we investigate how LLMs encode language identity by tracing the trajectory of representations across layers. We show that the transition from a English-aligned representation space to target-language token spaces is governed by a small and sparse set of language-specific dimensions that are consistent across layers. We introduc a simple, training-free method to identify and manipulate these dimensions, using as few as 50 sentences. Our inference-time interventions demonstrate that manipulating these dimensions enables controlled language switching while preserving semantic content, thereby revealing their interpretability and practical utility. Experimental results on multilingual generation control task confirm the effectiveness of our approach: it achieves strong performance in both monolingual and parallel settings, surpassing Neuron-Kojima by up to 12.69 points and Neuron-Tang by up to 5.87 points. Moreover, the parallel setting is generally stronger than the monolingual setting. These findings establish language-specific dimensions as a consistent and controllable mechanism underlying cross-lingual transfer in LLMs, opening new directions for both interpretability and efficient multilingual control.
@@ 7 Limitations
Our empirical evaluation is limited to multilingual generation control with a specific prompt type. While our method surpasses neuron-based baselines in this setting, the extent to which these gains transfer to downstream applications such as question answering and summarization remains open. A more comprehensive assessment would combine task-specific metrics with human judgments and examine robustness across prompts.
@@ A.1 Choosing an anchor layer for Llama2-13B
Figure 8: Choosing an anchor layer for monolingual setting. Ablation on Llama2–13B in the monolingual setting.

### v2

@@ Abstract
Large language models exhibit strong multilingual capabilities despite limited exposure to non-English data. Prior studies show that English-centric large language models map multilingual content into English-aligned representations at intermediate layers and then project them back into target-language token spaces in the final layer. From this observation, we hypothesize that this cross-lingual transition is governed by a small and sparse set of dimensions, which occur at consistent indices across the intermediate to final layers. Building on this insight, we introduce a simple, training-free method to identify and manipulate these dimensions, requiring only as few as 50 sentences of either parallel or monolingual data. Experiments on a multilingual generation control task reveal the interpretability of these dimensions, demonstrating that the interventions in these dimensions can switch the output language while preserving semantic content, and that it surpasses the performance of prior neuron-based approaches at a substantially lower cost. Our code is publicly available. [^1] https://github.com/ku-nlp/language-specific-dimensions
@@
Chengzhi Zhongthanks: This denotes equal contribution.1,   Fei Chengfootnotemark: 1,    Qianying Liu2, Yugo Murawaki1,Chenhui Chu1,Sadao Kurohashi1,2 1 Kyoto University, Japan    2 NII, Japan zhong@nlp.ist.i.kyoto-u.ac.jp {feicheng, murawaki, chu, kuro}@i.kyoto-u.ac.jp ying@nii.ac.jp
@@ 4.2 Data for Identification
4.3 Multilingual Generation Control
We evaluate with a pipeline of three metrics:
  • BLEU: evaluates translation quality for successful samples only.
4.4 MLQA
We add MLQA Lewis et al. (2020) as an additional task. MLQA is an extractive question answering benchmark, where the model is required to answer a question by copying a short span from the provided context. In this work, we extract 440 parallel examples from the MLQA dataset covering German, Spanish, and Chinese.
During evaluation, we use the following fixed English QA prompt format:
Answer the question using a short phrase directly copied from the context. Do not explain or generate additional text. Context: {context} Question: {question} Answer:
Although the prompt is always in English, we apply different intervention methods to steer the model to generate answers in a target language. The generated answers are evaluated against gold answers in the corresponding target language. We report performance using the F1 and Exact Match (EM) metrics.
5 Results and Discussion
We first search for an appropriate top-$K$ threshold for selecting language-specific dimensions. Therefore, we run the multilingual generation control experiment on Llama2-7B, applying an intervention at layer 19 with $\alpha\!=\!0.4$ in the parallel setting while varying $K$. As shown in Fig. 3, as $K$ increases, ACC improves but BLEU drops. The composite ACC*BLEU improves from 13.62 at K=50K\!=\!50, to 14.91 at K=100K\!=\!100 (+1.29), and to 15.80 at K=400K\!=\!400 (+2.18). Considering this trade-off, the most essential language-specific dimensions appear to be captured at around K=100K\!=\!100, which yields a smaller loss in semantic fidelity while still improving accuracy. However, to maximize performance, we set K=400K\!=\!400 for the subsequent experiments, which is about 7.8% of the hidden size for Llama2-13B and 9.8% for the other models.
In the monolingual setting, we need to choose an intermediate reference layer to compare with the final layer. We therefore set K=400K\!=\!400, apply an intervention at layer 19 with $\alpha\!=\!0.4$, and evaluate Llama2-7B on the multilingual generation control task while varying the layer $l$ from which dimensions are identified. As shown in Fig. 4, ACC and BLEU remain stable over a broad range of $l$, and the composite ACC*BLEU attains its best value near l=20l\!=\!20 but degrades at very deep layers. This indicates that the results are not highly sensitive to the choice of intermediate layer, and we use l=20l\!=\!20 in subsequent experiments.
@@ 5.1 Pre-examination on language-specific dimension identification
Llama2-7b Neuron Kojima 42.0 15.3 55.7 15.1 81.3 10.2 57.5 7.5 77.0 16.6 60.8 12.5 7.56
Neuron Tang 83.7 28.2 72.7 21.1 86.7 17.9 20.5 14.4 98.0 20.8 72.3 21.8 15.80
@@
Llama2-13b Neuron Kojima 15.7 7.3 30.3 9.5 96.0 15.9 64.5 8.6 14.0 4.3 47.4 12.2 5.80
Neuron Tang 83.7 32.5 14.7 23.5 99.3 19.1 64.5 8.6 42.0 25.8 63.7 22.4 14.23
@@
Llama3.1-8b Ours (Monolingual) 97.2 28.4 97.6 19.0 93.9 21.0 80.8 11.3 99.0 18.4 93.9 20.7 19.47
@@
Aya23-8b Ours (Monolingual) 68.0 21.7 69.4 14.7 22.8 20.3 69.8 10.8 58.7 14.0 55.6 16.5 9.34
@@
[…]
Output $\alpha\!=\!0.5$ 为什么一个人可能会选择在小型本地商店购物,而不是大型连锁商店?尽管价格更高(even if the prices are higher)
$\alpha\!=\!0.7$ 为什么一个人可能会选择在小型本地商店购物,而不是大型连锁商店,尽管价格更高?
$\alpha\!=\!0.9$ 为什么一个人可能会选择在小型本地商店购物,而不是大型链式(chain)商店,尽管价格(higher)高?
@@ 5.2 Main Results of Controlling Multilingual Generation
To assess generality, we evaluate on two latest models—Llama3.1–8B and Aya23–8B. On Llama3.1–8B in the monolingual setting, the overall ACC*BLEU reaches 18.70, the best among the settings we evaluated. The relatively lower performance of Llama3.1-8B under the parallel setting may be attributed to the use of hyperparameters tuned for Llama2-7B. On Aya23–8B in the parallel setting, it reaches 10.51. These results support our hypothesis about language-specific dimensions and indicate broad applicability across model families. We omit neuron-based baselines, as they are not compatible with these newer models. Besides BLEU, we additionally report BERTScore as an alternative semantic evaluation metric; the results are presented in Appendix A.4.
@@ 5.3 Layerwise Analysis of Language-Specific Dimensions
Model Method De Zh Es Overall
F1 EM F1 EM F1 EM F1 EM
Llama2-7b Neuron Tang 30.8 18.1 6.8 6.8 32.2 16.4 23.3 13.8
Ours (Monolingual) 27.4 15.7 8.1 5.9 25.9 13.4 20.5 11.7
Ours (Parallel) 25.4 14.5 7.3 5.5 26.5 13.6 19.8 11.2
Llama2-13b Neuron Tang 30.2 18.6 10.4 8.9 27.4 14.8 22.7 14.1
Ours (Monolingual) 33.1 22.7 10.8 9.3 34.7 17.0 26.2 16.4
Ours (Parallel) 33.9 20.2 11.7 10.0 36.7 18.6 27.4 16.3
Table 4: Results on MLQA. For reference, the English QA baseline without any intervention achieves F1/EM of 51.1/30.7 on Llama2-7B and 68.2/48.4 on Llama2-13B. All reported results use English prompts with interventions and are evaluated against gold answers in the target language.
5.4 Results of MLQA
We further evaluate the proposed intervention methods on MLQA, which constitutes a more challenging setting that requires the model to correctly follow instructions, comprehend the given context, and accurately locate the gold answer span.
The results are summarized in Table 4. Overall, our method achieves performance comparable to the neuron-based approach. While the neuron-based method performs slightly better on the 7B model, our approach outperforms it on the 13B model. This trend may be related to the increased stability of internal representations in larger models. However, all intervention-based methods perform substantially worse than the base model without intervention.
Beyond the effect of the intervention itself, this degradation also arises from the extractive nature of MLQA, where the model must generate target-language answers that do not appear in the English context. As a result, the task effectively shifts from extractive QA to cross-lingual answer generation, which inherently increases difficulty.
Nevertheless, both neuron-level and hidden-state-level intervention methods still demonstrate a non-trivial ability to induce language switching. This suggests that for more complex tasks such as QA, the model’s internal behavior becomes considerably more intricate: representation-level manipulations can influence the output language, albeit with a trade-off in overall task performance.
@@ 6 Conclusion
In this work, we investigate how LLMs encode language identity by tracing the trajectory of representations across layers. We show that the transition from a English-aligned representation space to target-language token spaces is governed by a small and sparse set of language-specific dimensions that are consistent across layers. We introduce a simple, training-free method to identify and manipulate these dimensions, using as few as 50 sentences. Our inference-time interventions demonstrate that manipulating these dimensions enables controlled language switching while preserving semantic content, thereby revealing their interpretability and practical utility. Experimental results on multilingual generation control task confirm the effectiveness of our approach: it achieves strong performance in both monolingual and parallel settings, surpassing Neuron-Kojima by up to 12.69 points and Neuron-Tang by up to 5.87 points. Moreover, the parallel setting is generally stronger than the monolingual setting. These findings establish language-specific dimensions as a consistent and controllable mechanism underlying cross-lingual transfer in LLMs, opening new directions for both interpretability and efficient multilingual control.
@@ 7 Limitations
Our empirical evaluation primarily focuses on multilingual generation control under a specific prompt design, with a limited study on question answering. While the results demonstrate the effectiveness of our method in controlled settings, its impact on a broader range of downstream tasks, such as mathematics and summarization, remains to be systematically examined. Future work should extend the evaluation scope to diverse tasks, incorporate task-specific metrics and human judgments, and assess robustness across different prompt formulations.
Acknowledgments
This work was supported by the “R&D Hub Aimed at Ensuring Transparency and Reliability of Generative AI Models” project of the Ministry of Education, Culture, Sports, Science and Technology, and by JSPS KAKENHI Grant Number JP23K28144, and by the Cross-ministerial Strategic Innovation Promotion Program (SIP) on “Integrated Health Care System” Grant No. JPJ012425.
@@ Appendix A Details of Intervention Setting
Figure 7: Choosing an anchor layer for monolingual setting. Ablation on Llama2–13B.
@@ A.2 Extra Experiments
A.3 Are 50 Sentences Sufficient for Identification
Model Method Fr De Zh Ja Es Overall
ACC BERT ACC BERT ACC BERT ACC BERT ACC BERT ACC BERT A*B
Llama2-7b Neuron Tang 83.7 0.945 72.7 0.941 86.7 0.918 20.5 0.922 98.0 0.945 72.3 0.935 67.58
Ours (Monolingual) 98.3 0.952 97.3 0.939 84.7 0.912 76.5 0.904 98.7 0.940 91.1 0.929 84.63
Ours (Parallel) 99.1 0.940 98.8 0.938 88.3 0.916 76.5 0.905 99.3 0.940 92.6 0.930 86.02
Llama2-13b Neuron Tang 83.7 0.948 14.7 0.942 99.3 0.925 64.5 0.896 42.0 0.949 63.7 0.936 59.62
Ours (Monolingual) 96.2 0.936 99.2 0.935 97.6 0.909 91.5 0.907 97.0 0.907 96.9 0.921 89.27
Ours (Parallel) 93.1 0.940 98.6 0.934 99.1 0.914 92.5 0.903 98.7 0.916 96.3 0.924 88.96
Table 5: Results of multilingual generation control with semantic preservation evaluated by XLM-R.
A.4 Results of BERTScore
In addition to BLEU, we also report semantic similarity using BERTScore with FacebookAI/xlm-roberta-large. As shown in Table 5, all methods exhibit comparable semantic preservation, with substantially smaller performance gaps than those observed with BLEU. Since BERTScore is evaluated only on samples where the generated output is in the correct target language, the improvement in the ACC*BERTScore metric mainly reflects the higher language accuracy achieved by our method, resulting in a clear advantage over the neuron-based baseline.